The speed of the slower motorcyclist is____mph
The speed of the faster motorcyclist is____mph.
Help please.

Answer by rajagopalan(174)

Two motorcyclists start at the same point
travel in opposite directions.
Speed of first moto cyclist=x
Speed of faster moto cyclist=x+4
combined speed=x+x+4=2x+4
distance in 3 hours=(2x+4)*3
6x+12=438
6x=438-12
6x=426
x=426/6=71
speed of first cyclist=71 mph
Speed of other=x+4=75 mph